OpenStudy (satan312):

Which is an equation in point-slope form for the given point and slope? Point: ; Slope a y+4=-5/6(x+8) b y-4=-5/6(x+8) c y+4=-5/6(x-8) d y-4=-5/6(x-8)

OpenStudy (anonymous):

I don't see the point or slope

OpenStudy (satan312):

ok one sec

OpenStudy (satan312):

point:(8-4) slope: -5/6

OpenStudy (anonymous):

Your answer is C

OpenStudy (satan312):

can you explain

OpenStudy (anonymous):

the standard point slope form is y-y1=m(x-x1) You input the 8 for x1 the -4 for y1 and the -5/6 for m and you get y-(-4)=(-5/6)(x-(8)) which simplifies to y+4=-5/6(x-8

OpenStudy (satan312):

ok can you help with a few more

OpenStudy (anonymous):

sure

OpenStudy (satan312):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2?a y+3=2(x+5) b y-3=2(x+5) c y+3=2(x-5) d y-3=2(x-5)

OpenStudy (anonymous):

D

OpenStudy (satan312):

explain?

OpenStudy (anonymous):

same thing put in 5 for y1 and 3 for x1 then put in 2 for m y-5=2(x-3) i just realized that all of the answer choices are wrong

OpenStudy (satan312):

oh ok thank you

OpenStudy (anonymous):

no problem